# Rainbow chains and numerical renormalisation group for accurate chiral conformal spectra

### v1

@@ 1 Introduction
Over the past decades, studying entanglement properties of quantum many-body states have generated many new insights in condensed-matter theory. In this quest, the notion of the entanglement spectrum [1] has proven particularly useful: It is obtained by interpreting the reduced density matrix of a subsystem $A$ as the thermal density matrix of an entanglement Hamiltonian $H_{\mathrm{ent},A}$, defined through
@@
where $e^{-\lambda_{\alpha}}$ and $|\alpha_{A}\rangle$ are the Schmidt values and vectors; the entanglement spectrum is conventionally defined as the set $\{\lambda_{\alpha}\}$. The “low-energy” structure of the entanglement spectrum (that is, the structure of leading Schmidt values) contains fingerprints of exotic physics, which would be challenging to diagnose with conventional tools. For example, the entanglement spectrum of a symmetry-protected topological (SPT) state contains degeneracies that reflect the symmetry group that protect the phase [2, 3], while the entanglement spectrum of two-dimensional fractional quantum Hall [1] and other chiral topologically ordered [4] states reflect the spectrum of the chiral conformal field theories (CFTs) that describe their edge spectra. The latter, combined with momentum-resolved entanglement spectra from MPS [5] or PEPS [6] simulations on finite cylinders, has become the method of choice to pinpoint chiral spin-liquid phases in numerics [7, 8, 9, 10, 11].
@@
In this paper, we propose an alternative approach, effectively reversing the relationship between entanglement spectra and BCFTs. Namely, we construct one-dimensional Hamiltonians whose entanglement spectra match those of a BCFT on a segment of length $\sim L$ with uniformly distributed sites, thereby much improving their finite-size scaling. We find that Hamiltonian terms in the bulk of such a chain decay exponentially from its middle, similar to the so-called rainbow chains [19, 20, 21, 22], which have been proposed as counterexamples to the entanglement area law in one dimension. While the volume-law scaling of entanglement entropy should generally be concerning for MPS simulations, we find that high-quality MPS representations can be built even for long free-fermion chains, allowing us to recover entanglement spectra in much more detail than it would be possible in computationally viable uniform chains.
For interacting theories, we also design a numerical renormalisation-group (NRG) scheme, based on Wilson’s NRG for the Kondo problem [23, 24], to obtain the ground state and low-lying excited states of these rainbow chains, a highly challenging task for variational optimisation methods like DMRG. We obtain these eigenstates in a nonstandard MPS form tailored to their entanglement structure, obviating potential difficulties due to volume-law entanglement. We also show how to convert this representation to a standard MPS form, thereby obtaining their entanglement spectra, which again recover CFT predictions to excellent accuracy.
@@
2 Rainbow-deformed chains
While the entanglement Hamiltonian (1) does not necessarily match the Hamiltonian of the system, they are closely related at one-dimensional conformal critical points [14]. The reduced density matrix of the quantum field theory is represented by a path integral with a branch cut across the subsystem in question: For 2D CFTs, the flexibility of 2D conformal transformations often allows this geometry to be mapped onto a rectangle, that is, the thermal density matrix of a finite segment of the CFT. That is, the entanglement Hamiltonian is equivalent to the Hamiltonian of the CFT itself, restricted to a segment, with some boundary conditions. In contrast to CFTs on closed manifolds, the spectrum of such a boundary CFT (BCFT) is typically generated by a single set of Virasoro generators, so its conformal blocks match those of the chiral version of the underlying CFT [16].
@@ 2 Rainbow-deformed chains
(d) CFT reduced density matrices in periodic boundary conditions can also be mapped onto thermal density matrices, using the conformal transformation (4). (Time runs radially.) The preimage of sites distributed uniformly in $w$-space between ±$\pm w_{2}/2$ remains symmetric under reflection across the imaginary axis. Reflection eigenvalues of Schmidt states can be used to distinguish descendant states with even and odd momentum.
In the particular case of the ground state of the finite open segment, $\operatorname{Re}z\in(-\pi/4,\pi/4)$, with an entanglement cut at the origin [Fig. 1(a)], such a mapping is given by [14]
@@
The mapping must be regularised by excluding a small circle $|z|<r$ from the domain of the path integral; discretising the theory to a chain of $L$ sites automatically achieves this, with a radius $r\sim 1/L$. It follows that the entanglement spectrum matches that of the BCFT on a segment of length $w_{0}\approx-\log r\approx\log L$, at inverse temperature $\beta=2\pi$: this leads to finite-size effects that scale with $\log L$, as observed in Refs. [13, 15]. Obtaining detailed and accurate information about the underlying CFT with this method would, therefore, require impractically long chains.
Instead, we propose discretising the continuum CFT to sites that are uniformly distributed after the conformal transformation [Fig. 1(b)]. That is, we place lattice sites at $w_{n}=-n\Delta$ ($n=0,1,\dots,L/2-1$), or rather, at $z_{\pm n}=\pm\arctan(e^{-n\Delta})$. Based on the discussion above, we expect the entanglement spectrum of the ground state of this system to capture the spectrum of the BCFT on a segment of length $w_{1}\sim L\Delta/2$, leading to much more favourable finite-size scaling.
To account for the uneven distribution of sites in the original position space $z$, we need to use a chain with unequal coupling strengths. In particular, Hamiltonian terms living on bonds have to represent the total Hamiltonian $\int\!\mathcal{H}(x)\,\mathrm{d}x\approx\mathcal{H}(x_{\mathrm{mid}})\,\ell_{\mathrm{bond}}$ on the same bond: While the bond length is proportional to the Jacobian $|\mathrm{d}z/\mathrm{d}w|$ of the conformal transformation, $\mathcal{H}$ is a component of the stress–energy tensor, and so scales as $|\mathrm{d}z/\mathrm{d}w|^{-2}$. Therefore, the coupling term as a whole scales inversely proportional to the scale factor:
@@
That is, the coupling strength is the highest in the middle of the chain, and decays exponentially towards the edges. We choose to evaluate (5) at the midpoint $w_{i+1/2}=(i+1/2)\Delta$ of the bond after the conformal transformation. In models with both single- and two-site terms, we also scale the former using (3) evaluated at the position of the site.
This prescription, however, does not apply directly to the bond across the entanglement cut, which is not mapped onto the real axis by the conformal mapping. For purely two-site Hamiltonians, we empirically found that evaluating (3) at the position of the last site for this bond leads to the best convergence. [^1] For Hamiltonians with both single- and two-site terms, we obtain the coupling strength of the central bond by evaluating (3) at $-(w_{1}+\Delta/4)$. The motivation for this prescription is perhaps the clearest in the Majorana/transverse-field-Ising chain studied in Sec. 3.2. In the Majorana form (8c), all couplings appear to live on “bonds” centred on former bond midpoints as well as sites. The length of these “bonds” is $\Delta/2$, so the left endpoint of the last one would be $\Delta/4$ to the left of the last site. The analogy with the rainbow chain persists, as the full set of coefficients on both bonds and sites still has the same structure, cf. Fig. 1(c). In the limit of large $w_{1}\sim L\Delta/2$, the middle section of the resulting chain matches the rainbow chain [19, 20, 21, 22] [Fig. 1(c)], a tight-binding chain with exponentially decaying couplings, studied as a counterexample of the area law of entanglement in local Hamiltonians [17, 25]. In fact, the NRG algorithm in Sec. 4 requires us to use such a chain, with uniform exponential decay all the way to the end, instead of the conformally deformed chain (3); however, this does not appear to change the behaviour of the Hamiltonian significantly.
2.1 Periodic boundary conditions: Reflection symmetry as a proxy for momentum
@@
almost identical to (3). We also set the interaction strength on the two bonds at the entanglement cut the same way as explained above.
@@ 3 Free-fermion examples
We first illustrate the ideas outlined above through the entanglement spectra of the ground states of quadratic fermionic Hamiltonians. These can be computed using only the spectrum of the single-particle Green’s function [26], which can be obtained at modest computational cost even for long chains with coupling strengths spanning several orders of magnitude. Furthermore, as every Schmidt vector of a Slater-determinant (Pfaffian) state is itself a Slater determinant (Pfaffian) [27, 28, 29, 30], accurate MPS representations of these wave functions can also be built without the need for computationally expensive variational optimisation (e.g., using DMRG) [31, 30, 32, 33].
In particular, we consider the nearest-neighbour tight-binding chain (the closest analogue of the rainbow chain) and the Kitaev chain at the topological transition, the latter of which is equivalent to a nearest-neighbour quadratic Hamiltonian of Majorana fermions. Upon the Jordan–Wigner transformation
$\sigma_{i}^{+}$ $=(-1)^{\sum^{i-1}n_{j}}c^{\dagger}_{i};$ $\sigma_{i}^{-}$ $=(-1)^{\sum^{i-1}n_{j}}c_{i};$ $\sigma^{z}$ $=2c^{\dagger}_{i}c_{i}-1,$ (6)
these are also equivalent to a nearest-neighbour XY chain and the critical transverse-field Ising (TFI) chain, respectively.
@@ 3.1 Complex fermions (XY chain)
where the hopping terms $J_{i+1/2}$ are given by Eq. (3). For uniform $J$, such a tight-binding chain is conformally critical, with one left- and one right-moving set of fermionic modes near the Fermi energy. The corresponding BCFT spectrum therefore matches that of one copy of the chiral fermion CFT, whose levels are $1,1,2,3,5,7,11,15,22,\dots$-fold degenerate in each charge sector, while the different charge sectors are offset by $(\Delta q)^{2}/2$, where $\Delta q$ is the charge deviation from half-filling [34].
Figure 2: (a–b) Ground-state entanglement spectra of the free-fermion chain (7) with the rainbow deformation (3) for $\Delta=1/4$ and chain length (a) $L=256$, (b) $L=258$. Up to 11 degenerate subspaces with the expected multiplicities $1,1,2,3,5,7,11,\dots$ are clearly seen in each charge sector. The red parabolas indicate the expected $(\Delta q)^{2}/2$ offset of entanglement energy between charge sectors.
(c) Entanglement spectrum of the rainbow-deformed (5) free-fermion ring for $\Delta=1/4$ and $L=256$. Colours indicate mirror-symmetry eigenvalues $+1$ (blue) and $-1$ (red) of the Schmidt vectors, which distinguish subspaces with even and odd momentum.
(e–f) Entanglement spectrum in the half-filled sector for a number of rainbow-deformed chains with (e) fixed $\Delta=1/4$, (f) fixed $L\Delta=32$ [highlighted in blue in panel (e)].
(g) Finite-size scaling of the deviation of the $k=6$ [11-fold degenerate, grey rectangles in panels (e,f)] multiplet as a function of the effective system size $w$ (cf. Appendix A). Both the average distance from the ideal entanglement energy (6 after scaling) and the standard deviation of the multiplet scales as $1/w^{2}$, with only weak dependence on $\Delta$ [indicated by colours, cf. panel (f)]. Straight lines are added as a guide to the eye; to reduce clutter, individual levels are not shown.
@@
We first obtained the ground-state entanglement spectrum of chains of length 256 and 258 with $\Delta=1/4$, shown in Fig. 2(a,b). As expected, the BCFT spectrum is realised in both cases; the different structure of the two plots is due to half-filling occurring at 64 and 64.5 fermions in each half of the chain, respectively. The conformal spectra are extremely accurate: the multiplicity of even the 11th conformal multiplet can easily be read off from the plot, in sharp contrast to a uniform chain of the same length [Fig. 2(d)], where only the first few levels of the conformal tower can be resolved.
@@
We also computed the entanglement spectrum for the rainbow-deformed ring shown in Fig. 1(d), together with the reflection eigenvalues of the Schmidt states. This is shown in Fig. 2(c): we again obtain very sharp conformal multiplets, which are also distinguished by their alternating reflection eigenvalues in each charge sector, as predicted in Sec. 2.1.
@@
To analyse the dependence of finite-size corrections on the parameters of the chains, we obtained entanglement spectra for a range of system sizes ($10\leq L\Delta\leq 64$ for $\Delta=1/2$, $1/4$, $1/8$, $1/16$, $1/32$). The length of the chain at fixed $\Delta$ has a strong effect on the width of conformal multiplets: they are much more clearly resolved for long chains [Fig. 2(e)]. By contrast, changing $\Delta$ even by a factor of 16 at fixed $L\Delta$ has only a modest effect on the entanglement spectrum [Fig. 2(f)]. This two-parameter scaling is shown in more detail in Fig. 2(g) on the example of the mean and standard deviation of the 7th (11-fold degenerate) conformal multiplet. For sufficiently long chains, both of these measures of finite-size effects scale as $1/w^{2}$, where $w$ is the effective system size after the conformal deformation, labelled $w_{1,2}$ in Fig. 1 (see also Appendix A), with little dependence on $\Delta$. This is consistent with earlier results on the scaling properties of the entanglement spectrum in the rainbow chain [20, 21], as well as the energy spectrum of CFT critical points on a lattice [35]. Similar to these works, we also find that the overall scale of entanglement gaps scales with system size as $1/w$ (not shown), consistent with volume-law scaling of entanglement entropy.

### v2

@@ 1 Introduction
Over the past decades, studying entanglement properties of quantum many-body states has generated many new insights in condensed-matter theory. In this quest, the notion of the entanglement spectrum [1] has proven particularly useful: It is obtained by interpreting the reduced density matrix of a subsystem $A$ as the thermal density matrix of an entanglement Hamiltonian $H_{\mathrm{ent},A}$, defined through
@@
where $e^{-\lambda_{\alpha}}$ and $|\alpha_{A}\rangle$ are the Schmidt values and vectors; the entanglement spectrum is conventionally defined as the set $\{\lambda_{\alpha}\}$. The “low-energy” structure of the entanglement spectrum (that is, the structure of leading Schmidt values) contains fingerprints of exotic physics, which would be challenging to diagnose with conventional tools. For example, the entanglement spectrum of a symmetry-protected topological (SPT) state contains degeneracies that reflect the symmetry group that protects the phase [2, 3], while the entanglement spectrum of two-dimensional fractional quantum Hall [1] and other chiral topologically ordered [4] states reflect the spectrum of the chiral conformal field theories (CFTs) that describe their edge spectra. The latter, combined with momentum-resolved entanglement spectra from MPS [5] or PEPS [6] simulations on finite cylinders, has become the method of choice to pinpoint chiral spin-liquid phases in numerics [7, 8, 9, 10, 11].
@@
In this paper, we propose an alternative approach, effectively reversing the relationship between entanglement spectra and BCFTs. Namely, we construct one-dimensional Hamiltonians whose entanglement spectra match those of a BCFT on a segment of length $\sim L$ with uniformly distributed sites, thereby much improving their finite-size scaling. We find that Hamiltonian terms in the bulk of such a chain decay exponentially from its centre, similar to the so-called rainbow chains [19, 20, 21, 22], which have been proposed as counterexamples to the entanglement area law in one dimension. While the volume-law scaling of entanglement entropy should generally be concerning for MPS simulations, we find that high-quality MPS representations can be built even for long free-fermion chains, allowing us to recover entanglement spectra in much more detail than it would be possible in computationally viable uniform chains.
For interacting theories, we also design a numerical renormalisation-group (NRG) scheme, based on Wilson’s NRG for the Kondo problem [23, 24], to obtain the ground state and low-lying excited states of these rainbow chains, a highly challenging task for variational optimisation methods like DMRG. We obtain these eigenstates in a nonstandard MPS form tailored to their entanglement structure, obviating potential difficulties due to volume-law entanglement. We also show how to convert this representation to a standard MPS, thereby obtaining their entanglement spectra, which again recover CFT predictions to excellent accuracy.
@@
2 Conformal transformations and rainbow chains
While the entanglement Hamiltonian (1) does not necessarily match the Hamiltonian of the system, they are closely related at one-dimensional conformal critical points [14]. The reduced density matrix of a quantum field theory is represented by a path integral with a branch cut across the subsystem in question: For 2D CFTs, the flexibility of 2D conformal transformations often allows this geometry to be mapped onto a rectangle, that is, the thermal density matrix of a finite segment of the CFT. That is, the entanglement Hamiltonian is equivalent to the Hamiltonian of the CFT itself, restricted to a segment with some boundary conditions. In contrast to CFTs on closed manifolds, the spectrum of such a boundary CFT (BCFT) is typically generated by a single set of Virasoro generators, so its conformal blocks match those of the chiral version of the underlying CFT [16].
@@ 2 Conformal transformations and rainbow chains
(d) CFT reduced density matrices in periodic boundary conditions can also be mapped onto thermal density matrices, using the conformal transformation (5). (Time runs radially in $z$-space.) The preimage of sites distributed uniformly in $w$-space between ±$\pm w_{2}/2$ remains symmetric under reflection across the imaginary axis. Reflection eigenvalues of Schmidt states can be used to distinguish descendant states with even and odd momentum.
In the particular case of the ground state of the finite open segment $\operatorname{Re}z\in(-\pi/4,\pi/4)$ with an entanglement cut at the origin [Fig. 1(a)], such a mapping is given by [14]
@@
The mapping must be regularised by excluding a small circle $|z|<r$ from the domain of the path integral; discretising the theory to a chain of $L$ sites automatically achieves this, with a radius $r\sim 1/L$. It follows that the entanglement spectrum matches that of the BCFT on a segment of length $w_{0}\simeq-\log r\simeq\log L$, at inverse temperature $\beta=2\pi$: this leads to finite-size effects that scale with $\log L$, as observed in Refs. [13, 15]. Obtaining detailed and accurate information about the underlying CFT with this method would, therefore, require impractically long chains.
Instead, we propose discretising the continuum CFT to sites that are uniformly distributed after the conformal transformation [Fig. 1(b)]. That is, we place lattice sites at $w_{n}=-n\Delta$ ($n=0,1,\dots,L/2-1$) or, rather, at $z_{\pm n}=\pm\arctan(e^{-n\Delta})$. Based on the discussion above, we expect the entanglement spectrum of the ground state of this system to capture the spectrum of the BCFT on a segment of length $w_{1}\sim L\Delta/2$, leading to much more favourable finite-size scaling.
To account for the uneven distribution of sites in the original position space $z$, we need to use a chain with unequal coupling strengths. Namely, Hamiltonian terms living on bonds have to represent the total Hamiltonian $\int\!\mathcal{H}(x)\,\mathrm{d}x\approx\mathcal{H}(x_{\mathrm{mid}})\,\ell_{\mathrm{bond}}$ on the same bond: While the bond length is proportional to the Jacobian $|\mathrm{d}z/\mathrm{d}w|$ of the conformal transformation, $\mathcal{H}$ is a component of the stress–energy tensor, and so scales as $|\mathrm{d}z/\mathrm{d}w|^{-2}$. Therefore, the coupling term as a whole is inversely proportional to the scale factor:
@@
That is, the coupling strength is the highest in the middle of the chain, and decays exponentially towards the edges. We choose to evaluate (6) at the midpoint $w_{i+1/2}=(i+1/2)\Delta$ of the bond after the conformal transformation.
This prescription, however, does not apply directly to the bond across the entanglement cut, which is not mapped onto the real axis by the conformal mapping. Empirically, we found that evaluating (3) at the position of the last site for this bond leads to the best convergence. In the limit of large $w_{1}\sim L\Delta/2$, the middle section of the resulting chain matches the rainbow chain [19, 20, 21, 22] [Fig. 1(c)], a tight-binding chain with exponentially decaying couplings
$J_{i}$ $=e^{w_{i}}\propto\begin{cases}1&\text{central bond}\\ \alpha^{2n-1}&\text{$n$ bonds away from centre;}\end{cases}$ $\alpha$ $=e^{-\Delta/2},$ (4)
studied as a counterexample of the area law of entanglement in local Hamiltonians [17, 25]. In fact, the NRG algorithm in Sec. 4 requires us to use such a chain, with uniform exponential decay all the way to the end, instead of the conformally deformed chain (3); however, this does not appear to change the behaviour of the Hamiltonian significantly.
Hamiltonians with on-site terms.
Figure 2: In a Hamiltonian with both bond (red) and on-site (blue) terms, both represent the Hamiltonian of the continuum CFT over a segment of length $\Delta/2$, half the original bond length. In particular, the two terms become identical in form for the transverse-field Ising model rewritten in terms of Majorana modes $\gamma_{i}$, Eq. (9c).
The construction so far is only concerned with purely two-site Hamiltonians. In models that also have on-site terms, we treat those too as the discretisation of a bond: Effectively, the post-mapping uniform chain has a doubled number of sites and bonds of length $\Delta/2$, centred on both former sites and bond midpoints (Fig. 2). This is particularly easily motivated for the transverse-field Ising chain studied in Sec. 3.2, which can be rewritten as a bond-only Hamiltonian (9c) of a doubled number of Majorana fermions. Since both on-site and interaction terms map to two-body terms, they are placed on equal footing by turning both sites and bonds of the original model into midpoints of Majorana bonds.
Therefore, terms on sites and non-central bonds are still scaled using (3, 4, 6), evaluated at the position of sites $w_{i}=i\Delta$ and bond midpoints $w_{i+1/2}=(i+1/2)\Delta$, respectively. To recreate the rainbow-chain prescription, however, the effective position of the bond across the entanglement cut becomes the leftmost site of the new effective chain, where the original leftmost site is the midpoint of a bond of length $\Delta/2$. That is, we have to evaluate (3, 4, 6) at $-\tilde{w}_{1}=-(w_{1}+\Delta/4)$; for the rainbow chain, the result is shown in Fig. 1(c).
Periodic boundary conditions: Reflection symmetry as a proxy for momentum.
@@
almost identical to (3). We also set the Hamiltonian terms of the two bonds across the entanglement cut the same way as explained above.
@@ 3 Free-fermion examples
We first illustrate the ideas outlined above through the ground-state entanglement spectra of quadratic fermionic Hamiltonians. These can be computed using only the spectrum of the single-particle Green’s function [26], which can be obtained at modest computational cost even for long chains with coupling strengths spanning several orders of magnitude. Furthermore, as every Schmidt vector of a Slater-determinant (Pfaffian) state is itself a Slater determinant (Pfaffian) [27, 28, 29, 30], accurate MPS representations of these wave functions can also be built without the need for computationally expensive variational optimisation (e.g., using DMRG) [31, 30, 32, 33].
In particular, we consider the nearest-neighbour tight-binding model and the Kitaev chain at the topological transition, the latter of which is equivalent to a nearest-neighbour quadratic Hamiltonian of Majorana fermions. Upon the Jordan–Wigner transformation
$\sigma_{i}^{+}$ $=(-1)^{\sum^{i-1}n_{j}}c^{\dagger}_{i},$ $\sigma_{i}^{-}$ $=(-1)^{\sum^{i-1}n_{j}}c_{i},$ $\sigma^{z}$ $=2c^{\dagger}_{i}c_{i}-1,$ (7)
these are also equivalent to a nearest-neighbour XY model and the critical transverse-field Ising (TFI) model, respectively. For both models, we consider three geometries:
  • We refer to the open chain with coupling terms given by (3) as the conformal chain, as it (approximately, cf. Appendix. A) maps onto a uniform chain under the conformal transformation (2) [Fig. 1(b)].
  • Likewise, we call the ring with coupling terms (6) the conformal ring [Fig. 1(d)].
  • We also consider rainbow chains (4) [Fig. 1(c)] whose Hamiltonian terms decay exponentially throughout.
@@ 3.1 Complex fermions (XY chain)
Figure 3: (a–b) Ground-state entanglement spectra of conformal (3) free-fermion chains (8) with $\Delta=1/4$ and length (a) $L=256$, (b) $L=258$. Up to 11 degenerate subspaces with the expected multiplicities $1,1,2,3,5,7,11,\dots$ are clearly seen in each charge sector. The red parabolas indicate the expected $(\Delta q)^{2}/2$ offset of entanglement energy between charge sectors.
(c) Entanglement spectrum of the conformal (6) free-fermion ring for $\Delta=1/4$ and $L=256$. Colours indicate mirror-symmetry eigenvalues $+1$ (blue) and $-1$ (red) of the Schmidt vectors, which distinguish subspaces with even and odd momentum.
(e–f) Entanglement spectrum in the half-filled sector for several conformal chains (6) with (e) fixed $\Delta=1/4$, (f) fixed $L\Delta=32$ [blue in panel (e)].
(g) Schmidt values in the $n=6$ [11-fold degenerate, grey background in panels (e, f)] conformal level as a function of the effective system size $L_{\mathrm{eff}}$ for conformal chain (top) and ring (bottom) geometries. $\Delta=1/4$, $1/2$ are omitted to reduce clutter. Fits to the form $1/(aL_{\mathrm{eff}}^{2}+bL_{\mathrm{eff}}^{3})$ are added as guides to the eye.
@@
where the hopping terms $J_{i+1/2}$ are given by one of Eqs. (3, 4, 6). For uniform $J$, such a tight-binding chain is conformally critical, with one left- and one right-moving set of fermionic modes near the Fermi energy. The corresponding BCFT spectrum therefore matches that of one copy of the chiral fermion (or, equivalently, chiral boson [34]) CFT, whose levels are $1,1,2,3,5,7,11,15,22,\dots$-fold degenerate in each charge sector, while the different charge sectors are offset by $(\Delta q)^{2}/2$, where $\Delta q$ is the charge deviation from half-filling [35].
@@
We first obtained the ground-state entanglement spectrum of conformal chains of length 256 and 258 with $\Delta=1/4$, shown in Fig. 3(a,b). As expected, the BCFT spectrum is realised in both cases; the different structure of the two plots is due to half-filling occurring at 64 and 64.5 fermions in each half of the chain, respectively. The conformal spectra are extremely accurate: The multiplicity of even the 11th conformal multiplet can easily be read off from the plot, in sharp contrast to a uniform chain of the same length [Fig. 3(d)], where only the first few levels of the conformal block can be resolved.
@@
We also computed the entanglement spectrum for a 256-site conformal ring, together with the reflection eigenvalues of the Schmidt states. This is shown in Fig. 3(c): we again obtain very sharp conformal multiplets, which are also distinguished by their alternating reflection eigenvalues in each charge sector, as predicted.
@@
To study the dependence of finite-size corrections on the parameters of the chains, we computed entanglement spectra for a range of system sizes ($8\leq L\Delta\leq 64$ for $\Delta=1/2$, $1/4$, $1/8$, $1/16$, $1/32$). We analyse these finite-size effects to the entanglement entropy and the structure of the entanglement spectrum in detail in Sec. 3.3. For now, we point out that the length of the chain at fixed $\Delta$ has a strong effect on the conformal multiplets: they are much more clearly resolved in long chains [Fig. 3(e)]. By contrast, changing $\Delta$ even by a factor of 16 at fixed $L\Delta$ has only a modest effect on the entanglement spectrum [Fig. 3(f)]. In fact, the two-parameter scaling can effectively be collapsed by introducing an effective chain length $L_{\mathrm{eff}}$ based on the von Neumann entropy (see Sec. 3.3 and Appendix A), as illustrated in Fig. 3(g). For the system sizes we can access, we find that the deviation of levels from the thermodynamic limit scales as $1/L_{\mathrm{eff}}^{2}$: This is consistent with earlier results on the scaling properties of the entanglement spectrum in the rainbow chain [20, 21], as well as the energy spectrum of CFT critical points on a lattice [36]. However, this scaling does not live up to closer scrutiny of the largest systems, see Sec. 3.3.